Paynesville Corporation manufactures and sells a preservative used in food and drug manufacturing. The company carries no inventories. The master budget calls for the company to manufacture and sell 122,000 liters at a budgeted price of $240 per liter this year. The standard direct cost sheet for one liter of the preservative follows.

Direct materials (2 pounds @ $15) $ 30
Direct labor (0.5 hours @ $46) 23


Variable overhead is applied based on direct labor hours. The variable overhead rate is $130 per direct-labor hour. The fixed overhead rate (at the master budget level of activity) is $65 per unit. All non-manufacturing costs are fixed and are budgeted at $2.3 million for the coming year.

At the end of the year, the costs analyst reported that the sales activity variance for the year was $732,000 unfavorable.

The following is the actual income statement (in thousands of dollars) for the year.

Sales revenue $ 28,138
Less variable costs
Direct materials 2,998
Direct labor 2,660
Variable overhead 7,130
Total variable costs $ 12,788
Contribution margin $ 15,350
Less fixed costs
Fixed manufacturing overhead 1,160
Non-manufacturing costs 1,340
Total fixed costs $ 2,500
Operating profit $ 12,850


During the year, the company purchased 198,000 pounds of material and employed 51,400 hours of direct labor.


Required:

a. Compute the direct material price and efficiency variances.
b. Compute the direct labor price and efficiency variances.
c. Compute the variable overhead price and efficiency variances.

(For all requirements, enter your answers in whole dollars. Indicate the effect of each variance by selecting "F" for favorable, or "U" for unfavorable. If there is no effect, do not select either option.)
Answer #1

Data for Material and Labor variance

Standard Actual
Qty Rate Amount Qty Rate Amount
Material 122,000 X 2 244,000 15.00 3,660,000 Material 198,000 15.141 2,998,000
Labour 122,000 X 0.5     61,000 46 2,806,000 Labour     51,400     51.75 2,660,000
Material Labor
Efficiency = (Actual Qty - Standard Qty) X Standard Rate =(198,000 - 244,000) X 15 $ (690,000) Favorable =(51,400 - 61,000) X 46 $ (441,600) Favorable
Rate Variance = (Actual Rate - Standard Rate) X Actual Qty =(15.141 - 15) X 198,000 $     28,000 Unfavorable =(51.75 - 46) X 51,400 $   295,600 Unfavorable
$ (662,000) Favorable $ (146,000) Favorable
Material price variance $                             28,000 U
Material quantity variance $                          (690,000) F
Labor rate variance $                           295,600 U
Labor efficiency variance $                          (441,600) F

For Variable Overhead

Variable Overhead
=(46,400 - 56,000) X 80 $(1,248,000) favorable - Efficiency
=(86.85 - 80) X 46,400 $     448,000 Unfavorable - Rate
$   (800,000) Favorable